(18.02) A survey of licensed drivers inquired about running red lights. One question asked, "Of every ten motorists who run a red light, about how many do you think will be caught?" The mean result for 911 respondents was x¯ = 1.77 and the standard deviation was s = 1.65. For this large sample, s will be close to the population standard deviation s, so suppose we know that s = 1.65.

Give a 95% confidence interval (±0.01) for the mean opinion in the population of all licensed drivers: ____ ± ____ .

The distribution of responses is skewed to the right rather than Normal. Will it strongly affect the z confidence interval for this sample?

No
Yes
Answer #1

Solution :

Given that,

Z/2 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (1.65  / 911)

= 0.11

Give a 95% confidence interval (±0.01) for the mean opinion in the population of all licensed drivers:

1.77 0.11

No